# The radius of statistical efficiency

### v1

@@ A.1 Sets of measures constrained by their support
Consider a linear subspace $\mathcal{K}\subseteq\mathcal{X}$ and a mean zero measure $\mu\in\mathcal{P}_{p}$. Recall that the compression $\Sigma_{\mu}|_{\mathcal{K}}$ of $\Sigma_{\mu}$ to $\mathcal{K}$ is the positive semidefinite quadratic form on $\mathcal{K}$ given by:
$\langle\Sigma_{\mu}v,v\rangle={\operatorname{\mathbb{E}}}_{x\sim\mu}[\langle x,v\rangle^{2}]\qquad\textrm{for all }v\in\mathcal{K}.$ (44)
@@
Lemma A.2 (Trace of covariance &\&& $W_{2}$ distance).
Consider a measure $\mu\in\mathcal{P}_{2}$ and let $\mathcal{K}$ be a proper subspace of $\mathbf{R}^{d}$. Then, the following equalities hold:
@@
${\operatorname{tr}}\,\Sigma_{\mu}|_{\mathcal{K}}=\mathop{\operatorname{\mathbb{E}}}_{x\sim\mu}{\operatorname{tr}}(P_{\mathcal{K}}xx^{\top}P_{\mathcal{K}})$ $=\mathop{\operatorname{\mathbb{E}}}_{x\sim\mu}\|P_{\mathcal{K}}x\|^{2}=\mathop{\operatorname{\mathbb{E}}}_{x\sim\mu}\operatorname{dist}^{2}(x,\mathcal{K}^{\perp})=W_{2}^{2}(\mu,(P_{\mathcal{K}^{\perp}})_{\#}\mu),$
@@
In light of (44), the matrix $\Sigma_{\mu}|_{\mathcal{K}}$ is singular if and only if the inclusion $\operatorname{supp}(\mu)\subset v^{\perp}$ holds for some $v\in\mathcal{K}\cap\mathbb{S}^{d-1}$. Define the set of measures for which $\Sigma_{\nu}|_{\mathcal{K}}$ is indeed singular:
$\mathcal{V}_{\mathcal{K}}=\big{\{}\nu\in\mathcal{P}_{2}\mid\operatorname{supp}(\nu)\subset v^{\perp}\text{~{}for some~{}}v\in\mathcal{K}\cap\mathbb{S}^{d-1}\big{\}}.$ (45)
The following theorem, the main result of the section, shows that the $W_{2}$ distance to $\mathcal{V}_{\mathcal{K}}$ is simply the minimal eigenvalue of $\Sigma_{\mu}|_{\mathcal{K}}$.
Theorem A.3: (Distance to $\mathcal{V}_{\mathcal{K}}$) Consider a measure $\mu\in\mathcal{P}_{2}$ and a proper linear subspace $\mathcal{K}$ of $\mathbf{R}^{d}$. Then, equality holds: $W_{2}^{2}(\mu,\mathcal{V}_{\mathcal{K}})=\lambda_{\min}(\Sigma_{\mu}|_{\mathcal{K}}),$ where $\mathcal{V}_{\mathcal{K}}$ is defined in (45). Moreover the distance of $\mu$ to $\mathcal{V}_{\mathcal{K}}$ is attained by the measure $(\mathsf{P}_{v^{\perp}})_{\#}\mu$, where $v\in\mathcal{K}$ is an eigenvector of $\Sigma_{\mu}|_{\mathcal{K}}$ corresponding to its minimal eigenvalue.
@@
For any vector $v\in\mathcal{K}\cap\mathbb{S}^{d-1}$, applying Lemma A.2 with ${\rm span}(v)$ in place of $\mathcal{K}$ yields:
$\langle\Sigma_{\mu}v,v\rangle={\operatorname{tr}}\,\Sigma_{\mu}|_{{\rm span}(v)}=W_{2}^{2}(\mu,(\mathsf{P}_{\mathcal{K}^{\perp}})_{\#}\mu).$ (46)
Let us now decompose $\mathcal{V}_{\mathcal{K}}$ into a union of simpler sets
$\mathcal{V}_{\mathcal{K}}=\bigcup_{v\in\mathcal{K}\cap\mathbb{S}^{d-1}}L_{v}\qquad\textrm{where}\qquad L_{v}:=\{\nu\in\mathcal{P}_{2}:\operatorname{supp}(\nu)\subset v^{\perp}\}.$
@@
$W_{2}(\mu,L_{v})\leq W_{2}(\mu,(\mathsf{P}_{{v}^{\perp}})_{\#}\mu)=\mathop{\operatorname{\mathbb{E}}}_{x\sim\mu}\operatorname{dist}(x,v^{\perp})^{2}\leq W_{2}(\mu,L_{v}).$
where the first inequality holds trivially, the equality follows from Lemma A.1, and the last inequality follows from (42) with $C=v^{\perp}$. Thus equality holds throughout. Using (46), we then conclude
$W_{2}(\mu,\mathcal{V}_{\mathcal{K}})=\inf_{v\in\mathcal{K}\cap\mathbb{S}^{d-1}}W_{2}(\mu,L_{v})=\inf_{v\in\mathcal{K}\cap\mathbb{S}^{d-1}}\langle\Sigma_{\mu}v,v\rangle=\lambda_{\min}(\Sigma_{\mu}|_{\mathcal{K}}),$
as claimed. It also follows immediately that for any minimal eigenvector ${v}$ of $\Sigma_{\mu}|_{\mathcal{K}}$, the pushforward measure $\nu=(\mathsf{P}_{{v}^{\perp}})_{\#}\mu$ attains the minimal $W_{2}$ distance of $\mu$ to $\mathcal{V}_{\mathcal{K}}$. ∎
@@ A.2 Spectral sets and functions of measures
In this section, we investigate a special class of functions on $\mathcal{P}_{2}$—called spectral—that depend on the measure only through the eigenvalues of its second moment matrix. This function class has close analogues in existing literature in matrix analysis and eigenvalue optimization. We postpone a detailed discussion on the related literature until the end of the section.
The following is the key definition.
@@
A function $F\colon\mathcal{P}_{2}\to\mathbf{R}\cup\{+\infty\}$ is called spectral if for any $\mu,\nu\in\mathcal{P}_{2}$ with the same second moment matrix $\lambda\left(\Sigma_{\mu}\right)=\lambda\left(\Sigma_{\nu}\right)$ equality $F(\mu)=F(\nu)$ holds.
A good example to keep in mind is the Schatten norm $F(\mu)=\|\lambda(\Sigma_{\mu})\|_{q}$ for any $q\in(0,\infty]$. Notice that in this example $F$ factors as a composition of the eigenvalue map $\lambda(\cdot)$ and the permutation-invariant function $f(v)=\|v\|_{q}$ on $\mathbf{R}^{d}$. Evidently, all spectral functions arise in this way.
@@
A function $f\colon\mathbf{R}^{d}_{+}\to\mathbf{R}\cup\{+\infty\}$ is called symmetric if equality $f(s(x))=f(x)$ holds for all $x\in\mathbf{R}^{d}_{+}$ and all permutations of coordinates $s(\cdot)$.
An elementary observation is that a function $F\colon\mathcal{P}_{2}\to\mathbf{R}\cup\{+\infty\}$ is spectral if and only if there exists a symmetric function $f:\mathbf{R}^{d}_{+}\to\mathbf{R}\cup\{+\infty\}$ satisfying
$F(\mu)=f(\lambda(\Sigma_{\mu}))\qquad\forall\mu\in\mathcal{P}_{2}.$
Concretely, the symmetric function $f$ can be obtained from $F$ by restricting to Gaussian measures with diagonal covariance $f(v)=F(N(0,\operatorname{Diag}(v)))$. The definition of spectral and symmetric functions easily extends to sets through indicator functions. Namely, a set $\mathcal{G}\subset\mathcal{P}_{2}$ is spectral if the indicator function $\delta_{\mathcal{G}}$ is spectral, while a set $G\subset\mathbf{R}^{d}_{+}$ is symmetric if the indicator $\delta_{G}$ is symmetric.
@@
$\mathcal{G}_{q}:=\{\mu\in\mathcal{P}_{2}:\lambda_{q}(\Sigma_{\mu})=\lambda_{q+1}(\Sigma_{\mu})\}.$ (47)
Although complicated, we may identify it with the symmetric set
$G_{q}=\{v\in\mathbf{R}^{d}_{+}:v^{(q)}=v^{(q+1)}\},$
where $v^{(i)}$ is the $i$’th largest coordinate of $v$. Indeed, equality $\mathcal{G}=\{\mu\in\mathcal{P}_{2}:\lambda(\Sigma_{\mu})\in G\}$ holds.
In this section, we will show that one can express the $W_{2}$-distance function to $\mathcal{G}$ purely in terms of the $\ell_{2}$-distance function to the much simpler set $G$. Indeed, we will prove the following theorem, which specialized to example (47) yields the expression
$W_{2}(\mu,\mathcal{G}_{q})=\operatorname{dist}_{2}\left(\sqrt{\lambda(\Sigma_{\mu})},G_{q}\right)=\tfrac{1}{\sqrt{2}}\left(\sqrt{\lambda_{q}}-\sqrt{\lambda_{q+1}}\right).$
Theorem A.6: (Distance to spectral sets in $\mathcal{P}_{2}$) Let $G\subset\mathbf{R}^{d}_{+}$ be a symmetric set. Define now the set of measures $\mathcal{G}=\{\nu\in\mathcal{P}_{2}:\lambda(\Sigma_{\nu})\in G\}.$ Then for any $\mu\in\mathcal{P}_{2}$, equality holds: $W_{2}(\mu,\mathcal{G})=\min_{v\in G}~{}\left\|\sqrt{\lambda(\Sigma_{\mu})}-\sqrt{v}\right\|_{2}.$
Indeed, we will prove a more general statement that applies to functions, with the distance replaced by the so-called Moreau envelope. We need some further notation to proceed. Let $(\mathcal{Y},\mathsf{d})$ be a metric space and consider a function $f\colon\mathcal{Y}\to\mathbf{R}\cup\{+\infty\}$. Then for any parameter $\rho>0$, the Moreau envelope and the proximal map of $f$ [47, 27], respectively, are defined as:
@@
$\operatorname{prox}_{\rho f}(y)$ $:=\operatornamewithlimits{argmin}_{y^{\prime}\in W}~{}f(y^{\prime})+\tfrac{1}{2\rho}\mathsf{d}^{2}(y,y^{\prime}).$
In particular, if $f$ is an indicator function of a set $Q$, then $f_{\rho}$ reduces to the squared distance function to $Q$, while $\operatorname{prox}_{\rho f}(w)$ is the nearest point projection.
@@
  • $(\mathcal{P}_{2},W_{2})$ The space $\mathcal{P}_{2}(\mathbf{R}^{d})$ equipped with the Wasserstein-2 distance $W_{2}(\cdot,\cdot)$.
@@
    where the square root is applied elementwise.
Notice that we are abusing notation by using the same symbol $W_{2}$ to denote the metric in all three spaces. The reason we are justified in doing so is that the three metric spaces are related by isometric embedding. Namely, the Wasserstein-2 distance between two Gaussian distributions $\mu=N(0,\Sigma_{\mu})$ and $\nu=N(0,\Sigma_{\nu})$ coincides with the Bures-Wasserstein distance between their covariance matrices:
@@
Theorem A.7: (Diagonal reduction) Consider a symmetric function $f\colon\mathbf{R}^{d}_{+}\to\mathbf{R}\cup\{+\infty\}$ and define the spectral function $F\colon\mathcal{P}_{2}\to\mathbf{R}\cup\{+\infty\}$ by setting $F(\mu)=f\left(\lambda\left(\Sigma_{\mu}\right)\right).$ Then equality holds: $F_{\rho}(\mu)=f_{\rho}(\lambda(\Sigma_{\mu}))\qquad\forall\mu\in\mathcal{P}_{2}(\mathbf{R}^{d}).$
@@
We begin with some notation. The symbol $O(d)$ will denote the set of orthogonal $d\times d$ matrices. The singular values for any matrix $A\in\mathbf{R}^{m\times n}$ (with $m\leq n$) in nonincreasing order we written as
$\sigma_{1}(A)\geq\sigma_{2}(A)\ldots\geq\sigma_{m}(A).$
We say that two matrices $A$ and $B$ admit a simultaneous ordered singular-value decomposition (SVD) if there exist matrices $U\in O(m),~{}V\in O(n)$ satisfying $U^{\top}AV=\operatorname{Diag}(\sigma(A))$ and $U^{\top}BV=\operatorname{Diag}(\sigma(B))$. The following trace inequality, essentially due to [64, 67], will play a central role in the section. The theorem as stated, along with a proof, may be found in [42, Theorem 4.6].
Theorem A.8 (von Neumann-Theobald).
@@
The following lemma shows that the Burer-Wasserstein distance can be written in terms of a Procrustes problem [7, Theorem 1].
@@
For any two matrices $A,B\in{\mathcal{S}}^{d}_{+}$ equality holds:
$W_{2}(A,B)=\min_{U\in O(d)}\|A^{1/2}-B^{1/2}U\|_{F}$
We will also need the following variational form of the Burer-Wasserstein distance; see for example [7, Section 3].
@@
For any two matrices $A,B\in\bm{S}^{d}_{+}$ equality holds:
$W_{2}^{2}(A,B)=\min_{x,y:~{}\operatorname{\mathbb{E}}[xx^{\top}]=A,\,\operatorname{\mathbb{E}}[yy^{\top}]=B}\operatorname{\mathbb{E}}\|x-y\|^{2}_{2}$
With these results in place, we are ready to start proving Theorem A.2. To this end, we will first establish the theorem in the Gaussian setting and then deduce the general case by a reduction. As the first step, we will estimate the $W_{2}$-distance of a matrix to an orbit of $B$ under conjugation:
$\mathcal{O}(B):=\{VBV^{\top}:V\in O(d)\}.$
Lemma A.11 (Distance to orbit: Gaussian case).
For any two matrices $A,B\in\bm{S}^{d}_{+}$ it holds:
$W_{2}(A,\mathcal{O}(B))=\left\|\sqrt{\lambda(A)}-\sqrt{\lambda(B)}\right\|_{2}.$ (48)
@@
$\{U\operatorname{Diag}(\lambda(B))U^{\top}:A=U\operatorname{Diag}(\lambda(A))U^{\top},~{}U\in O(d)\}.$ (49)
@@
To see the inequality $\leq$, consider an eigenvalue decomposition $A=U\operatorname{Diag}(\lambda(A))U^{\top}$ for some $U\in O(d)$. Then we have
$W_{2}(A,\mathcal{O}(B))\leq W_{2}(U\operatorname{Diag}(\lambda(A))U^{\top},U\operatorname{Diag}(\lambda(B))U^{\top})=W_{2}(\lambda(A),\lambda(B)),$
as claimed. Next, we show the reverse inequality $\geq$. To this end, set $\bar{A}:=A^{1/2}$ and $\bar{B}:=B^{1/2}$. Then for any $V\in O(d)$, we successively compute
$W_{2}(A,VBV^{\top})$ $=\inf_{U\in O(d)}\|\bar{A}-(V\bar{B}V^{\top})U\|_{F}^{2}$ (50)
$=\|\bar{A}\|^{2}+\|\bar{B}\|^{2}-2\sup_{U\in O(d)}\langle\bar{A},V\bar{B}(V^{\top}U)\rangle$ (51)
$=\|\bar{A}\|^{2}+\|\bar{B}\|^{2}-2\sup_{Z\in O(d)}\langle\bar{A},V\bar{B}Z^{\top}\rangle$ (52)
$\geq\|\bar{A}\|^{2}+\|\bar{B}\|^{2}-2\langle\sigma(\bar{A}),\sigma(\bar{B})\rangle$ (53)
$=\|\lambda(\bar{A})\|^{2}+\|\lambda(\bar{B})\|^{2}-2\langle\lambda(\bar{A}),\lambda(\bar{B})\rangle$ (54)

### v2

@@ A.1 Sets of measures constrained by their support
Consider a linear subspace $\mathcal{K}\subset\mathcal{X}$ and a measure $\mu\in\mathcal{P}_{2}$. Observe that the compression $\Sigma_{\mu}|_{\mathcal{K}}$ of $\Sigma_{\mu}$ to $\mathcal{K}$ is the positive semidefinite operator on $\mathcal{K}$ with quadratic form given by
$\langle\Sigma_{\mu}v,v\rangle=\mathop{\operatorname{\mathbb{E}}}_{x\sim\mu}\langle x,v\rangle^{2}\qquad\forall v\in\mathcal{K}.$ (47)
@@
Lemma A.2 (Trace of second moment &\& $W_{2}$ distance).
Consider a measure $\mu\in\mathcal{P}_{2}$ and let $\mathcal{K}$ be a linear subspace of $\mathbf{R}^{d}$. Then the following equalities hold:
@@
${\operatorname{tr}}\,\Sigma_{\mu}|_{\mathcal{K}}=\mathop{\operatorname{\mathbb{E}}}_{x\sim\mu}{\operatorname{tr}}(\mathsf{P}_{\mathcal{K}}xx^{\top}\mathsf{P}_{\mathcal{K}})$ $=\mathop{\operatorname{\mathbb{E}}}_{x\sim\mu}\|\mathsf{P}_{\mathcal{K}}x\|^{2}_{2}=\mathop{\operatorname{\mathbb{E}}}_{x\sim\mu}\operatorname{dist}(x,\mathcal{K}^{\perp})^{2}=W_{2}^{2}(\mu,(\mathsf{P}_{\mathcal{K}^{\perp}})_{\#}\mu),$
@@
In light of (47), the matrix $\Sigma_{\mu}|_{\mathcal{K}}$ is singular if and only if the inclusion $\operatorname{supp}(\mu)\subset v^{\perp}$ holds for some $v\in\mathcal{K}\cap\mathbb{S}^{d-1}$. Define the set of measures $\nu\in\mathcal{P}_{2}$ for which $\Sigma_{\nu}|_{\mathcal{K}}$ is singular:
$\mathcal{V}_{\mathcal{K}}:=\big\{\nu\in\mathcal{P}_{2}\mid\operatorname{supp}(\nu)\subset v^{\perp}\text{~for some~}v\in\mathcal{K}\cap\mathbb{S}^{d-1}\big\}.$ (48)
The following theorem—the main result of this section—shows that the squared $W_{2}$ distance of $\mu$ to $\mathcal{V}_{\mathcal{K}}$ is simply the minimal eigenvalue of $\Sigma_{\mu}|_{\mathcal{K}}$.
Theorem A.3: (Distance to $\mathcal{V}_{\mathcal{K}}$) Consider a measure $\mu\in\mathcal{P}_{2}$ and a nontrivial linear subspace $\mathcal{K}$ of $\mathbf{R}^{d}$. Then equality holds: $W_{2}^{2}(\mu,\mathcal{V}_{\mathcal{K}})=\lambda_{\min}(\Sigma_{\mu}|_{\mathcal{K}}),$ where $\mathcal{V}_{\mathcal{K}}$ is defined in (48). Moreover, the $W_{2}$ distance from $\mu$ to $\mathcal{V}_{\mathcal{K}}$ is attained by the measure $(\mathsf{P}_{v^{\perp}})_{\#}\mu$, where $v\in\mathcal{K}$ is any eigenvector of $\Sigma_{\mu}|_{\mathcal{K}}$ corresponding to its minimal eigenvalue.
@@
For any vector $v\in\mathcal{K}\cap\mathbb{S}^{d-1}$, applying Lemma A.2 with ${\rm span}(v)$ in place of $\mathcal{K}$ yields
$\langle\Sigma_{\mu}v,v\rangle={\operatorname{tr}}\,\Sigma_{\mu}|_{{\rm span}(v)}=W_{2}^{2}(\mu,(\mathsf{P}_{v^{\perp}})_{\#}\mu).$ (49)
Let us now decompose $\mathcal{V}_{\mathcal{K}}$ into a union of simpler sets:
$\mathcal{V}_{\mathcal{K}}=\bigcup_{v\in\mathcal{K}\cap\mathbb{S}^{d-1}}L_{v}\qquad\textrm{where}\qquad L_{v}:=\{\nu\in\mathcal{P}_{2}\mid\operatorname{supp}(\nu)\subset v^{\perp}\}.$
@@
$W^{2}_{2}(\mu,L_{v})\leq W^{2}_{2}(\mu,(\mathsf{P}_{{v}^{\perp}})_{\#}\mu)=\mathop{\operatorname{\mathbb{E}}}_{x\sim\mu}\operatorname{dist}(x,v^{\perp})^{2}\leq W^{2}_{2}(\mu,L_{v}),$
where the first inequality holds trivially, the equality follows from Lemma A.1, and the last inequality follows from (45) with $v^{\perp}$ in place of $\mathcal{K}$. Thus equality holds throughout. Using (49), we then conclude
$W^{2}_{2}(\mu,\mathcal{V}_{\mathcal{K}})=\inf_{v\in\mathcal{K}\cap\mathbb{S}^{d-1}}W^{2}_{2}(\mu,L_{v})=\inf_{v\in\mathcal{K}\cap\mathbb{S}^{d-1}}\langle\Sigma_{\mu}v,v\rangle=\lambda_{\min}(\Sigma_{\mu}|_{\mathcal{K}}),$
as claimed. It also follows immediately that for any minimal eigenvector v{v} of $\Sigma_{\mu}|_{\mathcal{K}}$, the pushforward measure $(\mathsf{P}_{{v}^{\perp}})_{\#}\mu$ attains the $W_{2}$ distance from $\mu$ to $\mathcal{V}_{\mathcal{K}}$. ∎
@@ A.2 Spectral sets and functions of measures
In this section, we investigate a special class of functions on $\mathcal{P}_{2}^{\circ}$—called spectral—that depend on the measure only through the eigenvalues of its covariance matrix. This function class has close analogues in existing literature on matrix analysis and eigenvalue optimization. We postpone a detailed discussion of the related literature until the end of the section.
For any matrix $A\in\bm{S}^{d}$, let $\lambda(A)=(\lambda_{1}(A),\ldots,\lambda_{d}(A))$ denote the vector of eigenvalues of $A$ arranged in nonincreasing order. The following is the key definition.
@@
A function $F\colon\mathcal{P}_{2}^{\circ}\to\mathbf{R}\cup\{+\infty\}$ is called spectral if for any $\mu,\nu\in\mathcal{P}_{2}^{\circ}$ satisfying $\lambda(\Sigma_{\mu})=\lambda(\Sigma_{\nu})$, the equality $F(\mu)=F(\nu)$ holds.
A good example to keep in mind is the Schatten $q$-norm $F(\mu)=\|\lambda(\Sigma_{\mu})\|_{q}$ for any $q\in[1,\infty]$ (where $\|\cdot\|_{q}$ denotes the $\ell_{q}$ norm on $\mathbf{R}^{d}$). Notice that in this example $F$ factors as a composition of the eigenvalue map $\lambda(\cdot)$ and the permutation-invariant function $f(v)=\|v\|_{q}$ on $\mathbf{R}^{d}$. Evidently, all spectral functions arise in this way.
@@
A function $f\colon\mathbf{R}^{d}_{+}\to\mathbf{R}\cup\{+\infty\}$ is called symmetric if the equality $f(s(x))=f(x)$ holds for all $x\in\mathbf{R}^{d}_{+}$ and all permutations of coordinates $s(\cdot)$.
An elementary observation is that a function $F\colon\mathcal{P}_{2}^{\circ}\to\mathbf{R}\cup\{+\infty\}$ is spectral if and only if there exists a symmetric function $f:\mathbf{R}^{d}_{+}\to\mathbf{R}\cup\{+\infty\}$ satisfying
$F(\mu)=f(\lambda(\Sigma_{\mu}))\qquad\forall\mu\in\mathcal{P}_{2}^{\circ}.$
Concretely, the symmetric function $f$ can be obtained from $F$ by restricting to Gaussian measures with diagonal covariance, i.e., $f(v)=F(\mathsf{N}(0,\operatorname{Diag}(v)))$. The definition of spectral and symmetric functions easily extends to sets through indicator functions. Namely, a set $\mathcal{G}\subset\mathcal{P}_{2}^{\circ}$ is spectral if the indicator function $\delta_{\mathcal{G}}$ is spectral, while a set $G\subset\mathbf{R}^{d}_{+}$ is symmetric if the indicator $\delta_{G}$ is symmetric.
@@
$\mathcal{G}_{q}:=\{\mu\in\mathcal{P}_{2}^{\circ}\mid\lambda_{q}(\Sigma_{\mu})=\lambda_{q+1}(\Sigma_{\mu})\}.$ (50)
Although complicated, we may represent $\mathcal{G}_{q}$ with the symmetric set
$G_{q}:=\{v\in\mathbf{R}^{d}_{+}\mid v^{(q)}=v^{(q+1)}\},$
where $v^{(i)}$ is the $i$’th largest coordinate of $v$. Indeed, equality $\mathcal{G}_{q}=\{\mu\in\mathcal{P}_{2}^{\circ}\mid\lambda(\Sigma_{\mu})\in G_{q}\}$ holds.
In this section, we will show that for any symmetric set $G\subset\mathbf{R}^{d}_{+}$, one can express the $W_{2}$ distance to the corresponding spectral set $\mathcal{G}=\{\nu\in\mathcal{P}_{2}^{\circ}\mid\lambda(\Sigma_{\nu})\in G\}$ purely in terms of the Hellinger distance to the much simpler set $G$. Indeed, we will prove the following theorem, which specialized to example (50) yields the expression
$W_{2}(\mu,\mathcal{G}_{q})={\operatorname{dist}}_{2}\Big(\sqrt{\lambda(\Sigma_{\mu})},G_{q}\Big)=\tfrac{1}{\sqrt{2}}\Big(\sqrt{\lambda_{q}}-\sqrt{\lambda_{q+1}}\Big).$
Theorem A.6: (Distance to spectral sets in $\mathcal{P}_{2}^{\circ}$) Let $G\subset\mathbf{R}^{d}_{+}$ be a symmetric set, and consider the corresponding spectral set of measures $\mathcal{G}:=\{\nu\in\mathcal{P}_{2}^{\circ}\mid\lambda(\Sigma_{\nu})\in G\}.$ Then for any $\mu\in\mathcal{P}_{2}^{\circ}$, equality holds: $W_{2}(\mu,\mathcal{G})=\min_{v\in G}\Big\|\sqrt{\lambda(\Sigma_{\mu})}-\sqrt{v}\Big\|_{2}.$
In fact, we will prove a more general statement that applies to functions, with the distance replaced by the so-called Moreau envelope. We need some further notation to proceed. Let $(\mathcal{Y},\mathsf{d})$ be a metric space and consider a function $f\colon\mathcal{Y}\to\mathbf{R}\cup\{+\infty\}$. Then for any parameter $\rho>0$, the Moreau envelope and the proximal map of $f$ [49, 27], respectively, are defined as:
@@
$\operatorname{prox}_{\rho f}(y)$ $:=\operatornamewithlimits{argmin}_{y^{\prime}\in\mathcal{Y}}~f(y^{\prime})+\tfrac{1}{2\rho}\mathsf{d}^{2}(y,y^{\prime}).$
In particular, if $f$ is an indicator function of a set $Q$, then $f_{\rho}$ is proportional to the squared distance function to $Q$, while $\operatorname{prox}_{\rho f}$ reduces to the nearest-point projection onto $Q$.
@@
  • $(\mathcal{P}_{2}^{\circ},W_{2})$ The space $\mathcal{P}_{2}^{\circ}(\mathbf{R}^{d})$ equipped with the Wasserstein-2 distance
    $W_{2}^{2}(\mu,\nu)=\min_{\pi\in\Pi(\mu,\nu)}\mathop{\operatorname{\mathbb{E}}}_{(x,y)\sim\pi}\|x-y\|_{2}^{2}.$
@@
    where the square root is applied componentwise.
Notice that we are abusing notation by using the same symbol $W_{2}$ to denote the metric in all three spaces. The reason we are justified in doing so is that the three metric spaces are related by isometric embedding. Namely, as shown in [51, Section 1.6.3], the Wasserstein-2 distance between two centered Gaussian distributions $\mu=\mathsf{N}(0,\Sigma_{\mu})$ and $\nu=\mathsf{N}(0,\Sigma_{\nu})$ coincides with the Bures-Wasserstein distance between their covariance matrices:
@@
Theorem A.7: (Diagonal reduction) Let $f\colon\mathbf{R}^{d}_{+}\to\mathbf{R}\cup\{+\infty\}$ be a symmetric function and $F\colon\mathcal{P}_{2}^{\circ}\to\mathbf{R}\cup\{+\infty\}$} be the spectral function given by $F(\mu)=f(\lambda(\Sigma_{\mu})).$ Then for any $\rho>0$ and $\mu\in\mathcal{P}_{2}^{\circ}$, equality holds: $F_{\rho}(\mu)=f_{\rho}(\lambda(\Sigma_{\mu})).$
@@
We begin with some notation. The symbol $O(d)$ will denote the set of real orthogonal $d\times d$ matrices. For any matrix $A\in\mathbf{R}^{m\times n}$, let $\sigma(A)=(\sigma_{1}(A),\ldots,\sigma_{m\wedge n}(A))$ denote the vector of singular values of $A$ arranged in nonincreasing order. We say that two matrices $A,B\in\mathbf{R}^{m\times n}$ admit a simultaneous ordered singular value decomposition (SVD) if there exist matrices $U\in O(m),~V\in O(n)$ satisfying $U^{\top}AV=\operatorname{Diag}(\sigma(A))$ and $U^{\top}BV=\operatorname{Diag}(\sigma(B))$. The following trace inequality, essentially due to [68, 71], will play a central role in the section. The result as stated, along with a proof, may be found in [41, Theorem 4.6].
Lemma A.8 (von Neumann-Theobald).
@@
The following lemma shows that the Bures-Wasserstein distance can be written in terms of a Procrustes problem [7, Theorem 1].
@@
For any two matrices $A,B\in\bm{S}^{d}_{+}$, equality holds:
$W_{2}(A,B)=\min_{U\in O(d)}\|A^{1/2}-B^{1/2}U\|_{F}.$
We will also need the following variational form of the Bures-Wasserstein distance; see for example [7, Section 3].
@@
For any two matrices $A,B\in\bm{S}^{d}_{+}$, equality holds:
$W_{2}^{2}(A,B)=\min_{\begin{subarray}{c}x,y:\,\operatorname{\mathbb{E}}[xx^{\top}]=A,\,\operatorname{\mathbb{E}}[yy^{\top}]=B\end{subarray}}\,\operatorname{\mathbb{E}}\|x-y\|^{2}_{2}.$
With these results in place, we are ready to start proving Theorem A.2. To this end, we will first establish the theorem in the Gaussian setting and then deduce the general case by a reduction. As the first step, we will compute the $W_{2}$ distance of a matrix to an orbit of $B$ under conjugation:
$\mathcal{O}(B):=\{VBV^{\top}\mid V\in O(d)\}.$
Lemma A.11 (Distance to orbit).
For any two matrices $A,B\in\bm{S}^{d}_{+}$, we have
$W_{2}(A,\mathcal{O}(B))=\Big\|\sqrt{\lambda(A)}-\sqrt{\lambda(B)}\Big\|_{2}=W_{2}(\lambda(A),\lambda(B)).$ (51)
@@
$\{U\operatorname{Diag}(\lambda(B))U^{\top}\mid A=U\operatorname{Diag}(\lambda(A))U^{\top},~U\in O(d)\}.$ (52)
@@
To see the inequality $\leq$ in (51), consider an eigenvalue decomposition $A=U\operatorname{Diag}(\lambda(A))U^{\top}$ for some $U\in O(d)$. Then we have
$W_{2}(A,\mathcal{O}(B))\leq W_{2}(U\operatorname{Diag}(\lambda(A))U^{\top},U\operatorname{Diag}(\lambda(B))U^{\top})=W_{2}(\lambda(A),\lambda(B)).$
Next, we show the reverse inequality. To this end, set $\bar{A}:=A^{1/2}$ and $\bar{B}:=B^{1/2}$. Then for any $V\in O(d)$, we successively compute
$W_{2}^{2}(A,VBV^{\top})$ $=\inf_{U\in O(d)}\|\bar{A}-(V\bar{B}V^{\top})U\|_{F}^{2}$ (53)
$=\|\bar{A}\|^{2}_{F}+\|\bar{B}\|^{2}_{F}-2\sup_{U\in O(d)}\langle\bar{A},V\bar{B}(V^{\top}U)\rangle$ (54)
$=\|\bar{A}\|^{2}_{F}+\|\bar{B}\|^{2}_{F}-2\sup_{Z\in O(d)}\langle\bar{A},V\bar{B}Z^{\top}\rangle$ (55)
$\geq\|\bar{A}\|^{2}_{F}+\|\bar{B}\|^{2}_{F}-2\langle\sigma(\bar{A}),\sigma(\bar{B})\rangle$ (56)
$=\|\lambda(\bar{A})\|^{2}_{2}+\|\lambda(\bar{B})\|^{2}_{2}-2\langle\lambda(\bar{A}),\lambda(\bar{B})\rangle$ (57)